# No evidence for $H_{0}$ anisotropy from Tully–Fisher or supernova distances

### v1

@@ 4 Results
We present results from applying our flow model to various data sets, focusing primarily on the CosmicFlows-4 Tully–Fisher relation W1 subsample. In Section˜4.1, we report constraints on the zero-point dipole magnitude and direction under different Galactic dust corrections and extinction priors, and assess the model preference based on Bayesian evidence. In Section˜4.2, we apply the same model to the 2MTF and SFI++ Tully–Fisher relation samples. We then examine the Pantheon+ supernova sample in Section˜4.3. In Table˜2 we summarise the recovered dipole magnitudes and Bayes factors for the CosmicFlows-4 W1 band, 2MTF, SFI++, and the Pantheon+ samples, while in Fig.˜4 we compare the inferred dipole direction with literature estimates, finding broad consistency. In Section˜4.4 we investigate the robustness of these results to assumptions about peculiar velocity corrections and the prior on the dipole amplitude. In Section˜4.5, we use mock Tully–Fisher relation catalogues to assess the detectability of a zero-point dipole as a function of survey sample size, evaluating the level of anisotropy that could be robustly inferred given realistic distance uncertainties. Finally, in Section˜4.6 we examine the bulk flow equivalent to the $H_{0}$ dipole, and compare it with both the bulk flow inferred when $\bm{V}_{\rm ext}$ is allowed a radial dependence and with $\Lambda$-cold dark matter expectation.
@@ 4.1 CosmicFlows-4 TFR W1 dipole
    Dipole in the zero-point with the extinction coefficient jointly inferred from the data, using one of three dust maps: Schlegel et al. (1998)Chiang (2023), or Planck Collaboration et al. (2016).
Figure 2: Comparison of the inferred Tully–Fisher relation zero-point dipole under different Galactic dust corrections using the CosmicFlows-4 Tully–Fisher relation W1 data: (a) fixed extinction from the CosmicFlows-4 catalogue which is based on  Schlegel et al. (1998), (b) E(BV)E(B\!-\!V) from Schlegel et al. (1998) with a sampled extinction coefficient $R_{\rm W1}$, (c) E(BV)E(B\!-\!V) from the Chiang (2023) map, and (d) E(BV)E(B\!-\!V) from the Planck Collaboration et al. (2016) map, both also with sampled $R_{\rm W1}$. The recovered zero-point dipole is unaffected by the extinction treatment and its magnitude is $\Delta_{\rm ZP}=0.087\pm 0.19$ mag. The values of redshift scatter $\sigma_{v}$ and Tully–Fisher relation scatter $\sigma_{\rm int}$ show no improvement over the isotropic model without a zero-point dipole, though the Bayesian evidence favours the anisotropic models with odds of 877:1877\!:\!1.
We find that neither the choice of dust map nor the sampling of the Galactic extinction coefficient has an appreciable effect on the inferred zero-point dipole. This indicates no significant degeneracy between the dipole and the Galactic extinction correction. We also consider two prior choices for the extinction coefficient $R_{\rm W1}$ (a narrow prior consistent with Yuan et al. (2013), and a broader, less informative alternative) and find that the results are similarly insensitive to this choice. All three models that include a dipole in the zero-point yield values of $\sigma_{v}$ and $\sigma_{\rm int}$ that are near-identical to those recovered by the isotropic model without a dipole. This would suggest that introducing a dipole does not appreciably improve the fit or capture additional trends in the data. However, upon considering a model comparison we find that the Bayesian evidence favours the anisotropic model with $\log\mathcal{B}=2.938$, or odds of 877:1877\!:\!1. We also find that the evidences of the models with different extinction treatments are nearly identical, indicating that the CosmicFlows-4 Tully–Fisher relation W1 data is insufficient to distinguish between the dust maps.
@@
Our conclusions are consistent with those of B25, who report a dipole magnitude of $0.063\pm 0.016$ mag, slightly lower than our inferred value and with a smaller quoted uncertainty. They find only a very weak support for the anisotropic model, reporting $\log\mathcal{B}=0.43$ in favour of the model including a zero-point dipole. However, they derive the Bayes factor using an approximation based on the Bayesian Information Criterion. We compare our results to B25 further in Section˜5.
@@ 4.3 Pantheon+ dipole
Furthermore, we find that the intrinsic supernova apparent magnitude scatter $\sigma_{\rm int}$ inferred by the model is $0.01\pm 0.01$ for Pantheon+ and $0.14\pm 0.01$ mag for Pantheon+Lane (both with and without the zero-point dipole). Brout et al. (2022) propagate an additional intrinsic scatter term directly into the covariance matrix, which explains why no further scatter is required (indeed the propagated scatter may be overestimated). Sah et al. (2024) argue that the Pantheon+ sample may suppress deviations from $\Lambda$-cold dark matter through this treatment, contrasting it with the covariance matrices of the Joint Light-Curve Analysis (JLA; Betoule et al. 2014) and Pantheon+Lane (Lane et al., 2024), which have a much smaller diagonal scatter. However, our inference of an additional intrinsic scatter of $0.14\pm 0.01$ mag indicates that the extra scatter propagated into the Pantheon+ covariance is justified.
@@
Sample Dipole Magnitude $\Delta H_{0}/H_{0}$ $\log\mathcal{B}$
CosmicFlows-4 TFR W1 $087\pm 0.019 0.0$$041\pm 0.009$ +2.938+2.938
2MTF $040\pm 0.024 0.0$$019\pm 0.011$ 0.381-0.381
SFI++ $051\pm 0.018 0.0$$024\pm 0.008$ +0.465+0.465
Pantheon+ $049\pm 0.013 0.0$$023\pm 0.006$
Pantheon+Lane $060\pm 0.013 0.0$$028\pm 0.006$
Table 2: Summary of inferred zero-point dipole magnitudes, fractional anisotropy in $H_{0}$, and Bayes factors in favour of the anisotropic model. Positive values of $\log\mathcal{B}$ indicate preference for the anisotropic model, negative values for the isotropic model. Only values with $|\log\mathcal{B}|\geq 2$ are considered as “decisive” evidence (Jeffreys, 1939).
Figure 4: Summary of the inferred zero-point dipole for each catalogue, given by its magnitude $\Delta_{\rm ZP}$ and direction in Galactic coordinates $(\ell,b)$, compared with literature estimates: the CosmicFlows-4 Tully–Fisher relation measurement (Boubel et al., 2025), the quasar dipole (Secrest et al., 2022b), the cluster scaling-relation dipole (Migkas et al., 2021), the CosmicFlows-4 bulk flow (Watkins et al., 2023), the Pantheon+ dipole (Sorrenti et al. 2023b, their $z<0.05$ sample), and the Local Group velocity in the cosmic microwave background frame (Planck Collaboration et al., 2020a). For the latter four, we plot the opposite direction to that reported, as our dipole is defined in the zero-point rather than in peculiar velocities. We show the inferred magnitude only for Boubel et al. (2025), as the remaining works measure dipoles in different quantities that we do not convert. All results are reported in the cosmic microwave background frame. Error bars indicate the 16th and 84th percentiles.
@@ 4.4 Model variations
Here we briefly consider how the results change on altering two aspects of the model. First, we consider the case in which the C15 reconstruction is not used to model the underlying density field and peculiar velocities. In this scenario, we omit the inhomogeneous Malmquist bias and model the velocity field with a constant flow $\bm{V}_{\rm ext}$. We then repeat the inference using the CosmicFlows-4 Tully–Fisher relation W1 subsample. In this case, the recovered dipole amplitude is $0.179\pm 0.023$ mag, significantly larger than the value obtained when accounting for peculiar velocities. The model selection in this “no peculiar velocity” scenario favours the inclusion of a zero-point dipole, effectively compensating for the unmodelled peculiar velocities, with a Bayes factor of $\log\mathcal{B}=6.73$. On the other hand, for Pantheon+ we only infer a significantly larger zero-point dipole uncertainty when not accounting for peculiar velocities, with the dipole magnitude being $0.051\pm 0.035$ mag. This highlights the importance of robustly accounting for peculiar velocities when investigating anisotropy, although it is important to bear in mind that these corrections may also absorb some anisotropy that may be present (as we discuss in Section˜1).
[…]
@@ 4.5 Dipole mock analysis
Above we reported a significant preference for a dipole in the CosmicFlows-4 Tully–Fisher relation W1 sample. We now inject a dipole of magnitude $0.087\,\mathrm{mag}$, corresponding to the posterior mean value inferred from the CosmicFlows-4 sample, and examine how the evidence ratio between anisotropic and isotropic models depends on sample size. The remaining injected parameters are listed in Table˜2, and an example redshift distribution of a mock sample is shown in Fig.˜1. Following Section˜3.4, we generate mock catalogues with $500$, $1000$, $2000$, $4000$, $8000$, $16000$, and $32000$ sources, drawing ten random realisations for each size. We then apply our flow model to the mock Tully–Fisher relation data, fitting both an isotropic model and one including a zero-point dipole, and compute the corresponding evidence ratio.
We present these results in Fig.˜6. The Bayes factor becomes decisive ($\log\mathcal{B}>2$ on the Jeffreys’ scale, Jeffreys 1939) in favour of the dipole model only at sample sizes of order $2000$. At smaller sizes it remains inconclusive and does not reliably discriminate between the models. Since the W1 sample after our quality cuts contains $3246$ galaxies, it is reassuring that the Bayes factor favours the zero-point dipole model near-decisively at that scale, assuming the injected dipole magnitude. The expected Bayes factor preference of mock samples of similar size to CosmicFlows-4 is likewise consistent with our results on the data. Fig.˜6 is also consistent with the absence of a significant dipole detection in the smaller 2MTF and SFI++ samples. For mock samples without an injected dipole the Bayes factor is always negative, favouring the isotropic model, although the strength of this preference increases only weakly with sample size. We note that these results are specific to a CosmicFlows-4 Tully–Fisher relation W1-like survey: while deeper surveys could in principle provide stronger constraints through larger samples, potential gains may be offset by increased absolute distance uncertainties at higher redshift.
Figure 6: Model comparison using CosmicFlows-4 Tully–Fisher relation W1-like mock data sets of varying size (see Section 3.4), testing an anisotropic zero-point dipole model against an isotropic alternative. Mocks are generated either with a zero-point dipole of magnitude $0.087$ (equal to the posterior mean of our constraints; red) or without it (blue). Larger values indicate stronger preference for the anisotropic model. The green lines indicate the sample size and Bayes factor of the CosmicFlows-4 Tully–Fisher relation W1 data, which are sufficient to yield “decisive” evidence on the Jeffreys’ scale ($\log\mathcal{B}>2$Jeffreys 1939). The shaded band shows the 16th–84th percentile range of the Bayes factor $\log\mathcal{B}$ across mock realisations of the same size.
@@ 4.6 Bulk flow measurement
The inferred dipole in $\bm{\Delta}_{\rm ZP}$ can be interpreted as a dipole in $H_{0}$ (see Eq. 9), or alternatively as a bulk flow linearly rising with distance under a cosmological model with isotropic expansion, which in turn corresponds to a bulk flow that increases linearly with distance. The bulk flow is typically defined as the average velocity within a volume $V$ of radius $R$ centred on the observer, assuming a top-hat filter (Watkins et al., 2009; Feldman et al., 2010; Nusser & Davis, 2011; Hoffman et al., 2015; Watkins & Feldman, 2015; Nusser, 2016; Scrimgeour et al., 2016; Feix et al., 2017; Hellwing et al., 2018; Peery et al., 2018; Whitford et al., 2023; Watkins et al., 2023). In our case, three possible contributions enter:
@@
where the first term is the velocity field of C15 scaled by $\beta^{\star}$, the second is the external bulk flow $\bm{V}_{\rm ext}$ (optionally position dependent), and the third is the flow implied by the zero-point dipole (or equivalently $H_{0}$ dipole), here approximated with a linear Hubble law (the full expression yields negligible differences), for which we assume $H_{0}=73.04~\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ (Riess et al., 2022). This is for illustration only; the exact results depend minimally on the assumed $H_{0}$ value. Sources along the zero-point dipole direction are intrinsically fainter (i.e. have a larger zero-point), inducing a dipole in their apparent magnitudes but not in their redshifts, since they are assumed to be uniformly distributed in volume. If the model accounts only for a velocity dipole and not the zero-point dipole, it assigns all sources the same zero-point. Consequently, sources along the dipole direction appear fainter and are inferred as more distant. To match the isotropy of redshifts, the model compensates by introducing a negative velocity dipole along the zero-point direction, hence the minus sign in Eq.˜22.
@@
Moreover, we compare the magnitude of this bulk flow to the $\Lambda$-cold dark matter expectation, computed with a top-hat filter of radius $R$ (see e.g. Sec. 6.1.2 of Boruah et al. 2020), assuming Planck 2018 cosmology (Planck Collaboration et al., 2020b). This comparison is essential: while the data may show a strong preference for a dipole model (or more generally the existence of a bulk flow), it could still be entirely consistent with $\Lambda$-cold dark matter.
@@
We show the resulting bulk flow in Fig.˜7 for three extensions of C15: (i) a constant $\bm{V}_{\rm ext}$ as in our main analysis, (ii) a constant $\bm{V}_{\rm ext}$ combined with $\bm{\Delta}_{\rm ZP}$, and (iii) a radially varying $\bm{V}_{\rm ext}(R)$. (In all cases the velocities contained within C15 are included.) As expected, the first case asymptotes to a constant above $R\approx 60~h^{-1}\,\mathrm{Mpc}$, where the constant $\bm{V}_{\rm ext}$ dominates over the decaying contribution from C15. The second case behaves similarly at small radii, since the zero-point dipole contribution scales with distance, but yields a larger bulk flow above $R\approx 40~h^{-1}\,\mathrm{Mpc}$ and then grows linearly once the dipole contribution dominates. Finally, the radially varying $\bm{V}_{\rm ext}(R)$ produces a bulk flow slightly lower than the constant $\bm{V}_{\rm ext}$ case at small radii, higher at $R\approx 40$$60~h^{-1}\,\mathrm{Mpc}$, and decaying again at larger distances.
@@
Figure 7: Bulk flow as a function of radius $R$ inferred from the CosmicFlows-4 W1 sample. We first show the bulk flow from Carrick et al. (2015) only and then introduce three extensions of it: (i) a constant external flow $\bm{V}_{\rm ext}$, (ii) a constant $\bm{V}_{\rm ext}$ combined with the zero-point dipole $\bm{\Delta}_{\rm ZP}$, and (iii) a radially varying $\bm{V}_{\rm ext}(R)$. For reference, the shaded band shows the $\Lambda$-cold dark matter expectation computed with a top-hat filter of radius $R$, and the top panel displays the distribution of redshift distances in the CosmicFlows-4 W1 sample. All bands show the range between the $16$th to 84th percentile. All cases are consistent with $\Lambda$-cold dark matter expectations except the zero-point dipole model at large radii, where the (extrapolated) linearly growing dipole dominates. This is however disfavoured relative to the more flexible radially varying $\bm{V}_{\rm ext}$ model. For $\bm{V}_{\rm ext}(\bm{r})$, we place spline knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, indicated by vertical dashed lines.
@@
In Fig.˜8 we show the magnitude of $\bm{V}_{\rm ext}(\bm{r})$ that is superimposed on the C15 velocity field. The magnitude rises from $200~\mathrm{km}\,\mathrm{s}^{-1}$ at small radii to about $300~\mathrm{km}\,\mathrm{s}^{-1}$ at $50~h^{-1}\,\mathrm{Mpc}$, after which it rapidly decays. $\bm{V}_{\rm ext}$ is interpolated between knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, with the magnitude at each knot sampled uniformly between 0 and $500~\mathrm{km}\,\mathrm{s}^{-1}$Fig.˜8 demonstrates that this more flexible model is inconsistent with a linearly growing bulk flow, which would signal $H_{0}$ anisotropy. Furthermore, above $80~h^{-1}\,\mathrm{Mpc}$ inferred $\bm{V}_{\rm ext}$ is nearly consistent with zero, while the uncertainty band broadens largely due to the lack of data and the posterior transitioning to the prior. We verify that these conclusions remain unchanged when using linear rather than cubic interpolation between the knots.

### v2

@@ 4 Results
We present results from applying our flow model to various data sets, focusing primarily on the CosmicFlows-4 Tully–Fisher relation W1 subsample. In Section˜4.1, we report constraints on the zero-point dipole magnitude and direction under different Galactic dust corrections and extinction priors, and assess the model preference based on Bayesian evidence. In Section˜4.2, we apply the same model to the 2MTF and SFI++ Tully–Fisher relation samples. We then examine the Pantheon+ supernova sample in Section˜4.3. In Table˜2 we summarise the recovered dipole, external flow, and Bayes factors for the CosmicFlows-4 W1 band, 2MTF, SFI++, and the Pantheon+ samples, while in Fig.˜4 we compare the inferred dipole direction with literature estimates, finding broad consistency. In Section˜4.4 we investigate the robustness of these results to assumptions about peculiar velocity corrections and the prior on the dipole amplitude. In Section˜4.5, we use mock Tully–Fisher relation catalogues to assess the detectability of a zero-point dipole as a function of survey sample size, evaluating the level of anisotropy that could be robustly inferred given realistic distance uncertainties. Finally, in Section˜4.6 we examine the bulk flow equivalent to the $H_{0}$ dipole, and compare it with both the bulk flow inferred when $\bm{V}_{\rm ext}$ is allowed a radial dependence and with $\Lambda$-cold dark matter expectation.
@@ 4.1 CosmicFlows-4 TFR W1 dipole
    Dipole in the zero-point with the extinction coefficient jointly inferred from the data, using one of three dust maps: Schlegel_1998Chiang_2023, or Planck_2016.
Figure 2: Comparison of the inferred Tully–Fisher relation zero-point dipole under different Galactic dust corrections using the CosmicFlows-4 Tully–Fisher relation W1 data: (a) fixed extinction from the CosmicFlows-4 catalogue which is based on  Schlegel_1998, (b) E(BV)E(B\!-\!V) from Schlegel_1998 with a sampled extinction coefficient $R_{\rm W1}$, (c) E(BV)E(B\!-\!V) from the Chiang_2023 map, and (d) E(BV)E(B\!-\!V) from the Planck_2016 map, both also with sampled $R_{\rm W1}$. The recovered zero-point dipole is unaffected by the extinction treatment and its magnitude is $\Delta_{\rm ZP}=0.087\pm 0.19$ mag. The values of redshift scatter $\sigma_{v}$ and Tully–Fisher relation scatter $\sigma_{\rm int}$ show no improvement over the isotropic model without a zero-point dipole, though the Bayesian evidence favours the anisotropic models with odds of 877:1877\!:\!1.
We find that neither the choice of dust map nor the sampling of the Galactic extinction coefficient has an appreciable effect on the inferred zero-point dipole. This indicates no significant degeneracy between the dipole and the Galactic extinction correction. We also consider two prior choices for the extinction coefficient $R_{\rm W1}$ (a narrow prior consistent with Yuan_2013, and a broader, less informative alternative) and find that the results are similarly insensitive to this choice. All three models that include a dipole in the zero-point yield values of $\sigma_{v}$ and $\sigma_{\rm int}$ that are near-identical to those recovered by the isotropic model without a dipole. This would suggest that introducing a dipole does not appreciably improve the fit or capture additional trends in the data. However, upon considering a model comparison we find that the Bayesian evidence favours the anisotropic model with $\log\mathcal{B}=2.938$, or odds of 877:1877\!:\!1. We also find that the evidences of the models with different extinction treatments are nearly identical, indicating that the CosmicFlows-4 Tully–Fisher relation W1 data is insufficient to distinguish between the dust maps.
@@
Our conclusions are consistent with those of Boubel_2025, who report a dipole magnitude of $0.063\pm 0.016$ mag, slightly lower than our inferred value and with a smaller quoted uncertainty. They find only a very weak support for the anisotropic model, reporting $\log\mathcal{B}=0.43$ in favour of the model including a zero-point dipole. However, they derive the Bayes factor using an approximation based on the Bayesian Information Criterion. We compare our results to Boubel_2025 further in Section˜5.
@@ 4.3 Pantheon+ dipole
Furthermore, we find that the intrinsic supernova apparent magnitude scatter $\sigma_{\rm int}$ inferred by the model is $0.01\pm 0.01$ for Pantheon+ and $0.14\pm 0.01$ mag for Pantheon+Lane (both with and without the zero-point dipole). Brout_2022 propagate an additional intrinsic scatter term directly into the covariance matrix, which explains why no further scatter is required (indeed the propagated scatter may be overestimated). Sah argue that the Pantheon+ sample may suppress deviations from $\Lambda$-cold dark matter through this treatment, contrasting it with the covariance matrices of the Joint Light-Curve Analysis (JLA; Betoule_2014) and Pantheon+Lane (Lane_2024), which have a much smaller diagonal scatter. However, our inference of an additional intrinsic scatter of $0.14\pm 0.01$ mag indicates that the extra scatter propagated into the Pantheon+ covariance is justified.
@@
Sample $\Delta_{\mathrm{ZP}}$ $\ell_{\Delta_{\mathrm{ZP}}}~[\mathrm{deg}]$ $b_{\Delta_{\mathrm{ZP}}}~[\mathrm{deg}]$ $V_{\mathrm{ext}}~[\mathrm{km}\,\mathrm{s}^{-1}]$ $\ell_{\mathrm{ext}}~[\mathrm{deg}]$ $b_{\mathrm{ext}}~[\mathrm{deg}]$ $\Delta H_{0}/H_{0}$ $\log\mathcal{B}$
CF4 W1 $0.087\pm 0.019$ $128\pm 12$ $9\pm 10$ $158\pm 20$ $303\pm 8$ $17\pm 5$ $0.041\pm 0.009$ +2.94+2.94
SFI++ $0.051\pm 0.018$ $84\pm 25$ $5\pm 17$ $168\pm 23$ $311\pm 7$ $4\pm 6$ $0.024\pm 0.008$ +0.47+0.47
2MTF $0.041\pm 0.023$ $130\pm 83$ $37\pm 29$ $150\pm 21$ $317\pm 9$ $\pm 7 0.0$$019\pm 0.011$ 0.38-0.38
Pantheon+ $0.049\pm 0.014$ $135\pm 56$ $49\pm 18$ $212\pm 39$ $312\pm 10$ $9\pm 7$ $0.023\pm 0.006$
Pantheon+Lane $0.060\pm 0.014$ $84\pm 50$ $52\pm 14$ $220\pm 34$ $312\pm 8$ $21\pm 6$ $0.028\pm 0.006$
Table 2: Summary of zero-point dipole and external flow parameters, as well as the fractional anisotropy in $H_{0}$, and Bayes factors in favour of the anisotropic model. Positive values of $\log\mathcal{B}$ indicate preference for the anisotropic model, negative values for the isotropic model. Only values with $|\log\mathcal{B}|\geq 2$ are considered as “decisive” evidence (Jeffreys).
Figure 4: Summary of the inferred zero-point dipole for each catalogue, given by its magnitude $\Delta_{\rm ZP}$ and direction in Galactic coordinates $(\ell,b)$, compared with literature estimates: the CosmicFlows-4 Tully–Fisher relation measurement (Boubel_2025), the quasar dipole (Secrest_2022), the cluster scaling-relation dipole (Migkas_2021), the CosmicFlows-4 bulk flow (Watkins_2023), the Pantheon+ dipole (Sorrenti_2023, their $z<0.05$ sample), and the Local Group velocity in the cosmic microwave background frame (Planck_2020). For the latter four, we plot the opposite direction to that reported, as our dipole is defined in the zero-point rather than in peculiar velocities. We show the inferred magnitude only for Boubel_2025, as the remaining works measure dipoles in different quantities that we do not convert. All results are reported in the cosmic microwave background frame. Error bars indicate the 16th and 84th percentiles.
@@ 4.4 Model variations
Here we briefly consider how the results change on altering two aspects of the model. First, we consider the case in which the Carrick_2015 reconstruction is not used to model the underlying density field and peculiar velocities. In this scenario, we omit the inhomogeneous Malmquist bias and model the velocity field with a constant flow $\bm{V}_{\rm ext}$. We then repeat the inference using the CosmicFlows-4 Tully–Fisher relation W1 subsample. In this case, the recovered dipole amplitude is $0.179\pm 0.023$ mag, significantly larger than the value obtained when accounting for peculiar velocities. The model selection in this “no peculiar velocity” scenario favours the inclusion of a zero-point dipole, effectively compensating for the unmodelled peculiar velocities, with a Bayes factor of $\log\mathcal{B}=6.73$. On the other hand, for Pantheon+ we only infer a significantly larger zero-point dipole uncertainty when not accounting for peculiar velocities, with the dipole magnitude being $0.051\pm 0.035$ mag. This highlights the importance of robustly accounting for peculiar velocities when investigating anisotropy, although it is important to bear in mind that these corrections may also absorb some anisotropy that may be present (as we discuss in Section˜1).
Second, we consider the impact of the assumed prior on the dipole amplitude. Previously we have taken this to be a uniform prior on the magnitude of the dipole vector; we now make this uniform on the Cartesian components (the choice made by Boubel_2025). This implicitly induces a prior on the dipole magnitude proportional to the square of the magnitude, pushing the posterior towards larger amplitudes. We repeat the analysis on the CosmicFlows-4 Tully–Fisher relation W1 subsample using this component-based prior, and compare the results to those obtained with a uniform prior on the magnitude in Fig.˜5. As expected, the recovered dipole magnitude is larger under the component prior, while the inferred direction is essentially unchanged. We find that the dipole magnitude is $0.095\pm 0.019$ mag with the flat-component prior, compared to $0.087\pm 0.019$ mag with the flat-magnitude prior, a $\sim 0.5\sigma$ difference.
@@ 4.5 Dipole mock analysis
Above we reported a significant preference for a dipole in the CosmicFlows-4 Tully–Fisher relation W1 sample. We now inject a dipole of magnitude $0.087\,\mathrm{mag}$, corresponding to the posterior mean value inferred from the CosmicFlows-4 sample, and examine how the evidence ratio between anisotropic and isotropic models depends on sample size. The remaining injected parameters are listed in Table˜1, and an example redshift distribution of a mock sample is shown in Fig.˜1. Following Section˜3.4, we generate mock catalogues with $500$, $1000$, $2000$, $4000$, $8000$, $16000$, and $32000$ sources, drawing ten random realisations for each size. We then apply our flow model to the mock Tully–Fisher relation data, fitting both an isotropic model and one including a zero-point dipole, and compute the corresponding evidence ratio.
We present these results in Fig.˜6. The Bayes factor becomes decisive ($\log\mathcal{B}>2$ on the Jeffreys’ scale, Jeffreys) in favour of the dipole model only at sample sizes of order $2000$. At smaller sizes it remains inconclusive and does not reliably discriminate between the models. Since the W1 sample after our quality cuts contains $3246$ galaxies, it is reassuring that the Bayes factor favours the zero-point dipole model near-decisively at that scale, assuming the injected dipole magnitude. The expected Bayes factor preference of mock samples of similar size to CosmicFlows-4 is likewise consistent with our results on the data. Fig.˜6 is also consistent with the absence of a significant dipole detection in the smaller 2MTF and SFI++ samples. For mock samples without an injected dipole the Bayes factor is always negative, favouring the isotropic model, although the strength of this preference increases only weakly with sample size. We note that these results are specific to a CosmicFlows-4 Tully–Fisher relation W1-like survey: while deeper surveys could in principle provide stronger constraints through larger samples, potential gains may be offset by increased absolute distance uncertainties at higher redshift.
Figure 6: Model comparison using CosmicFlows-4 Tully–Fisher relation W1-like mock data sets of varying size (see Section 3.4), testing an anisotropic zero-point dipole model against an isotropic alternative. Mocks are generated either with a zero-point dipole of magnitude $0.087$ (equal to the posterior mean of our constraints; red) or without it (blue). Larger values indicate stronger preference for the anisotropic model. The green lines indicate the sample size and Bayes factor of the CosmicFlows-4 Tully–Fisher relation W1 data, which are sufficient to yield “decisive” evidence on the Jeffreys’ scale ($\log\mathcal{B}>2$Jeffreys). The shaded band shows the 16th–84th percentile range of the Bayes factor $\log\mathcal{B}$ across mock realisations of the same size.
@@ 4.6 Bulk flow measurement
The inferred dipole in $\bm{\Delta}_{\rm ZP}$ can be interpreted as a dipole in $H_{0}$ (see Eq. 9), or alternatively as a bulk flow linearly rising with distance under a cosmological model with isotropic expansion, which in turn corresponds to a bulk flow that increases linearly with distance. The bulk flow is typically defined as the average velocity within a volume $V$ of radius $R$ centred on the observer, assuming a top-hat filter (Watkins_2009; Feldman_2010; Nusser_2011; Hoffman_2015; Watkins_2015; Nusser_2016; Scrimgeour_2016; Feix_2017; Hellwing_2018; Peery_2018; Whitford_2023; Watkins_2023). In our case, three possible contributions enter:
@@
where the first term is the velocity field of Carrick_2015 scaled by $\beta^{\star}$, the second is the external bulk flow $\bm{V}_{\rm ext}$ (optionally position dependent), and the third is the bulk-flow equivalent of the zero-point dipole (or equivalently $H_{0}$ dipole), here approximated with a linear Hubble law (the full expression yields negligible differences). We assume here $H_{0}=73.04~\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ (Riess_2022) for illustrative purposes: the exact results depend minimally on the assumed $H_{0}$. Sources along the zero-point dipole direction are intrinsically fainter (i.e. have a larger zero-point), inducing a dipole in their apparent magnitudes but not in their redshifts, since they are assumed to be uniformly distributed in volume. If the model accounts only for a velocity dipole and not the zero-point dipole, it assigns all sources the same zero-point and hence absolute magnitude on average. Consequently, sources along the dipole direction are inferred as more distant. To match the isotropy of redshifts, the model compensates by introducing a negative velocity dipole along the zero-point direction, hence the minus sign in Eq.˜22.
@@
We compare the magnitude of this bulk flow to the $\Lambda$-cold dark matter expectation, computed with a top-hat filter of radius $R$ (see e.g. Sec. 6.1.2 of Boruah_2019), assuming Planck 2018 cosmology (Planck_2020_cosmo). This comparison is essential: while the data may show a strong preference for a dipole model (or more generally the existence of a bulk flow), it could still be entirely consistent with $\Lambda$-cold dark matter.
@@
We show the resulting bulk flow in Fig.˜7 for three extensions of Carrick_2015: (i) a constant $\bm{V}_{\rm ext}$ as in our main analysis, (ii) a constant $\bm{V}_{\rm ext}$ combined with $\bm{\Delta}_{\rm ZP}$, and (iii) a radially varying $\bm{V}_{\rm ext}(R)$. (In all cases the velocities contained within Carrick_2015 are included.) As expected, the first case asymptotes to a constant above $R\approx 60~h^{-1}\,\mathrm{Mpc}$, where the constant $\bm{V}_{\rm ext}$ dominates over the decaying contribution from Carrick_2015. The second case behaves similarly at small radii, since the zero-point dipole contribution scales with distance, but yields a larger bulk flow above $R\approx 40~h^{-1}\,\mathrm{Mpc}$ and then grows linearly once the dipole contribution dominates. Finally, the radially varying $\bm{V}_{\rm ext}(R)$ produces a bulk flow slightly lower than the constant $\bm{V}_{\rm ext}$ case at small radii, higher at $R\approx 40$$60~h^{-1}\,\mathrm{Mpc}$, and decaying again at larger distances.
@@
Figure 7: Bulk flow as a function of radius $R$ inferred from the CosmicFlows-4 W1 sample. We first show the bulk flow from Carrick_2015 only and then introduce three extensions of it: (i) a constant external flow $\bm{V}_{\rm ext}$, (ii) a constant $\bm{V}_{\rm ext}$ combined with the zero-point dipole $\bm{\Delta}_{\rm ZP}$, and (iii) a radially varying $\bm{V}_{\rm ext}(R)$. For reference, the shaded band shows the $\Lambda$-cold dark matter expectation computed with a top-hat filter of radius $R$, and the top panel displays the distribution of redshift distances in the CosmicFlows-4 W1 sample. All bands show the range between the $16$th and 84th percentile. (The width of the Carrick_2015 is due to the $\beta^{\star}$ posterior.) All cases are consistent with $\Lambda$-cold dark matter expectations except the zero-point dipole model at large radii, where the (extrapolated) linearly growing dipole dominates. This is however disfavoured relative to the more flexible radially varying $\bm{V}_{\rm ext}$ model. For $\bm{V}_{\rm ext}(\bm{r})$, we place spline knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, indicated by vertical dashed lines.
@@
In Fig.˜8 we show the magnitude of $\bm{V}_{\rm ext}(\bm{r})$ that is superimposed on the Carrick_2015 velocity field. The magnitude rises from $200~\mathrm{km}\,\mathrm{s}^{-1}$ at small radii to about $300~\mathrm{km}\,\mathrm{s}^{-1}$ at $50~h^{-1}\,\mathrm{Mpc}$, after which it rapidly decays. $\bm{V}_{\rm ext}$ is interpolated between knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, with the magnitude at each knot sampled uniformly between 0 and $500~\mathrm{km}\,\mathrm{s}^{-1}$Fig.˜8 demonstrates that this more flexible model is inconsistent with a linearly growing bulk flow, which would signal $H_{0}$ anisotropy. Furthermore, above $80~h^{-1}\,\mathrm{Mpc}$ inferred $\bm{V}_{\rm ext}$ is nearly consistent with zero, while the uncertainty band broadens largely due to the lack of data and the posterior transitioning to the prior. The posterior on the magnitude does not appreciably broaden beyond $100~h^{-1}\,\mathrm{Mpc}$, as sufficient data remain to constrain it at that distance; beyond this, we assume a constant extrapolation. The $\bm{V}_{\rm ext}(\bm{r})$ model has higher evidence than the $\bm{V}_{\rm ext}+\bm{\Delta}_{\rm ZP}$ model, and we verify that these conclusions are robust to the interpolation scheme and number and locations of the knots.